OpenStudy (anonymous):

Solve 3 log6 2 + 2 log6 3 - log6 x = 2 Question 10 options: 1) x = 2 2) x = 6 3) x = 15 4) x = 36

OpenStudy (dumbcow):

x=2

OpenStudy (anonymous):

The trick for solving all types of logarithms are using and memorizing the properties. Therefore four properties we are going to use for this exercise is as follows:\[\log_{A}B+\log_AC=\log_AB*C\];\[\log_{A}B-\log_AC=\log_A(B/C)\] ; \[\log_AB ^{C}=C*\log_AB\] and finally \[A ^{\log_AB}=B\]. Now that you know the basic rules, lets apply them for this particular exercise: \[3\log_{6}2+2\log_63-\log_6x=2\] can be simplified using the third property as\[\log_62^{3}+\log_63^{2}-\log_6x=2\] which can simplify by the first property\[\log_6(2^{3}*3^{2})-\log_6x=2\] and by the second as\[\log_6{((8*9)/x)=2}\]. Using the fourth property we have that:\[6^{\log_{6}(72/x)}=6^{2}\] which gives us\[72/x = 36\] that simplifies into\[x=72/36=2\]. Peace, Love and Happiness from Puerto Rico